# Structured Packing in LLM Training Improves Long Context Utilization

### v6

@@ 3.2 Experimental Results
Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. We perform three trainings using different training datesets and report the mean and standard deviation.
@@
RoPe scale Training Method arXiv Code Code &
method data Haskell Python CUDA All arXiv
Naive C# SPLiCeBM25 6.25 (.08) 4.84 (.19) 3.55 (.11) 2.84 (.12) 3.72 (.13) 3.88 (.12)
@@
SPLiCe works with various context extension results Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension. FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. This allows for computationally efficient long-context fine-tuning. To show that our method works more generally, we have also evaluated it using more standard approaches. That is, we checked the effectiveness of the method when context extension is done in all layers, and parameters of Rotary Positional Encodings are either adjusted as in CodeLlama (Rozière et al., 2023), adjusted with YaRN (Peng et al., 2023), or left without changes. Table 3 shows the results.
3.3 Design Choices Analysis
Choice of k As already mentioned in the previous section, the choice of the retrieval method (BM25 or Contriver) plays a minor role. In Appendix C.1, we study the role of $k$. We first observe that its choice has a minor impact on the results, with $k=1$ being the best choice. We leave for further work a more detailed analysis of this hyperparameter.
Document ordering We discovered that the ordering of the documents collected by the retrieval procedure might be a subtle issue. We consider three choices of ORDER function in Algorithm 1: identity (do not permute the documents), reverse (reverse the order of documents) and a random shuffle. In Appendix C.1, we present results for $270$M models showing negligible differences. However, we observed differences for large models, elaborated in Section 4.2.4, in which random shuffling is the best choice for the CodeLlama context extension method.
Corpus chunking Performing retrieval from the whole corpus might be cumbersome due to its scale. However, we find that chunking the corpus into smaller parts still brings improvements. In particular, for the experiments with English Wikipedia and BM25 presented in Table 3, we used a manageable size of $2$GB.
4 Large-Scale Models
In this section, we show that SPLiCe can improve the long context performance of large-scale language models. To this end, we use $3$B and $7$B parameter models and tasks that test in-context learning (Section 4.2.1), question answering (Section 4.2.2), and in-context information retrieval capabilities (Section 4.2.3). We provide perplexity results in Appendix E.
We also emphasize that those improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B models were tuned on $2$B tokens. All of the models were pre-trained on $1$T tokens.
4.1 Experimental Setup
For $3$B experiments, we continue training OpenLLaMA $3$B v2 (Geng & Liu, 2023; Geng, 2023) on a 50/50 mixture of RedPajama prepared in the standard way and C prepared using SPLiCeBM25. For $7$B ones, we continue the training of OpenLLaMA $7$B v2 on 75/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25. StackExchange is part of RedPajama (TogetherComputer, 2023), and C data come from StarCoder (Li et al., 2023b).
We train with $32$K context length using $5.4$B tokens for $3$B models and $2$B for $7$B models (recall that OpenLLaMA models have the $2$K context length). We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023).
Regarding the context extensions method we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test six models:
[…]
where Baseline denotes the standard data preparation method. Hyperparameter details are located in Appendix A.
4.2 Experimental Results
We present perplexity measurements in Appendix E. In particular, following (Anthropic, 2023), we examine perplexity improvements in the function of the token position in the document. Our findings reveal that SPLiCe models demonstrate superior perplexity across all distances. Notably, the improvement in perplexity is more pronounced for tokens positioned later in the document, indicating enhanced language modeling capabilities.
For the remainder of this section, we focus on testing the in-context learning, question answering and information retrieval abilities on our models using downstream tasks.
4.2.1 In-Context Learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li & Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{190,380,780,1560\}$ (which correspond roughly to $\{4,8,16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on test set. In Table 4, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia, see Table 5, confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the domination of SPLiCe.
In Table 4 and Table 5, we report mean improvement $\Delta$ along with $95\%$ bootstrap confidence intervals. See Appendix I for related histograms.
Table 4: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). For TREC, we average across $50$ sets of in-context examples for $3$B models and $10$ sets for $7$B models. We use $\Delta$ [confidence interval] to denote mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I).
TREC
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (780) 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ]
$8$K (380) 66.5 75.8 9.3 [8.1,10.8]{8.1,10.8]}8.1 , 10.8 ]
$4$K (190) 65.1 72.8 7.7 [6.5,9.0]{6.5,9.0]}6.5 , 9.0 ]
[…]
$16$K (780) 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ]
$8$K (380) 72.1 76.2 4.1 [2.9,5.7]{2.9,5.7]}2.9 , 5.7 ]
$4$K (190) 66.9 69.7 2.8 [1.5,4.2]{1.5,4.2]}1.5 , 4.2 ]
[…]
$16$K (780) 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ]
$8$K (380) 76.6 76.5 -0.1 [1.0,1.1]{-1.0,1.1]}- 1.0 , 1.1 ]
$4$K (190) 68.1 69.1 1.0 [0.4,2.9]{-0.4,2.9]}- 0.4 , 2.9 ]
Table 5: Average performance on DBpedia (Lehmann et al., 2015). We average results across $40$ sets of in-context examples, with $\Delta$ [confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence interval. Due to the size of the evaluation dataset for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
DBpedia
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (190) 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
[…]
$16$K (190) 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
[…]
$16$K (190) 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
Table 6: Average performance on HotPotQA (Yang et al., 2018). We average results across $7$ sets of in-context examples, with with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
HotPotQA
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
4.2.2 Question Answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021; Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 7, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, see Table 6. We stress that in both cases, the results measure in-context capabilities as neither of the datasets was used during training.
Table 7: Two-shot F1 performance on the validation subset of Qasper (Dasigi et al., 2021). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
QASPER
Model
Context length
(#examples)
Baseline SPLiCe
[…]
4.2.3 Key Retrieval Performance
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models, see Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions. We provide the results for FoT models in Appendix D.
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
4.2.4 Shuffling
In our experiments with large $7$B CodeLlama models, we have noted slight improvements in handling long inputs when shuffling documents inside the context, that is, setting Order in the Algorithm 1 to be a random shuffle. We provide a comparison on TREC in Table 17.
5 Related Work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). A recent work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, (Han et al., 2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of this approach in large-scale language modeling, specifically with OpenLLaMA $3$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in (Han et al., 2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a similar method for preparing the training data. However, in our work, we focus on tuning the models for long context, up to $32$K tokens, whereas (Shi et al., 2023b) focuses on training the models from scratch, with context length up to $8$K tokens. In addition to that, we provide ablations regarding context extension methods.
6 Limitations and Future Work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work. Note that as BM25 is a syntactic retriever, it is not guaranteed that the documents returned by it are semantically related. Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
Scaling Our method requires further studies to understand its scaling properties beyond the range presented in the paper.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies of the design choices of SPLiCe with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, answers, etc.) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
7 Conclusions
In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves the perplexity of the language modeling task and performance on downstream tasks. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
8 Impact Statement
This paper presents work whose goal is to improve the long context utilization ability of language models through structuring training data. It bears standard risks associated with advancing the field of LLMs (such as misuse and generation of malicious/misleading content). For a detailed overview, we refer to (Bender et al., 2021; Weidinger et al., 2021).
Acknowledgments
Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al., 2023), and the architectural details can be found in Table 8. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al., 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al., 2021). For context extension, we use Focused Transformer (FoT)(Tworkowski et al., 2023), CodeLlama (Rozière et al., 2023) (CL) and YaRN (Peng et al., 2023). Table 9 presents the details about both standard and long-context pretraining.
Parameter/Model Size 270M 3B 7B
Vocab Size 32000 32000 32000
Embedding Size 1024 3200 4096
Num Attention Layers 12 26 32
Num Attention Heads 8 32 32
Head Size 128 100 128
MLP Hidden Size 4096 8640 11008
FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24]
Table 9: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B parameter OpenLLaMAv2 model (Geng, 2023). Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al., 2023) and no-FoT to other methods (Naive, YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023)).
Stage Parameter/Model Size 270M 3B 7B
Pretraining Context 2K 2K 2K
Tokens 6.3B 1T 1T
Long Context Pretraining Context 32KFoT, 16Kno-FoT 32K 32K
Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT
Start Learning Rate 5e-5 1.5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6 1.5e-6
Warmup Steps 250 1000 1000FoT, 200no-FoT
Tokens 1B 5.4B 2B
Appendix B Detailed Results For Medium-Size Models
In this section, we extend results presented in Section 3. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 18.

### v7

@@ 3.3 Detailed study with medium models
Table 5: Perplexity fine-tune on a $/50$ 50 data mixture of RedPajama and C code. We report the mean and standard deviation. Interestingly, training on the code data with SPLiCe improves general long-context performance on arXiv.
@@
3.4 Summary of the results
Structuring data improves performance. We observe (see Tables 15) that all tested variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe, see Tables 2 and 5, that the positive effects of SPLiCe transfer between domains; when training on code and evaluating on Qasper/arXiv and when training on web data and evaluating on code. This is a very interesting effect, which indicates that training induces generic long-context capabilities. Following that, we conjecture that better data mixtures and synthetic long-context data could bring further benefits.
SPLiCe improves the performance of various context extension methods SPLiCe only changes the input structure, and conjecturally, it should work with any context extension method. Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023). FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. In Table 9 in Appendix B.2, we confirm that SPLiCe integrates seamlessly with methods that extend the context in all layers. Specifically, we test adjusting Rotary Positional Encodings according to either the recipe as in CodeLlama (Rozière et al., 2023), or YaRN (Peng et al., 2023), or left without changes.
4 Related work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). Work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, Han et al. (2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase and long context performance, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of our approach in large-scale language modeling, specifically with OpenLLaMA $3$B, $7$B, and CodeLlama $13$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in Han et al. (2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a method for preparing the training data that closely resembles SPLiCe with default settings ($k=1$ and the identity as Order). The key differences are that (Shi et al., 2023b) focuses on training from scratch while we aim to achieve long context capabilities with short and cheap fine-tuning. Next, we use longer context lengths, up to $32$K tokens, whereas (Shi et al., 2023b) utilizes $8$K. Finally, we provide detailed analysis regarding the design choices, such as the benefits of data reordering in Section 3.2.4 and different values of $k$ in Appendix C.1.
5 Limitations and future work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method, SPLiCe, can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange questions and answers) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
6 Conclusions
In this work, we present SPLiCe, a method of constructing training examples for long-context language models. It utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves performance on downstream tasks and the language modeling abilities of LLMs. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
Acknowledgments and Disclosure of Funding
Appendix A Architecture
The architecture of our models is based on LLaMA [Touvron et al., 2023], and the architectural details can be found in Table 6. Briefly speaking, our architecture is similar to the one introduced in [Vaswani et al., 2017] with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding [Su et al., 2021]. For context extension, we use Focused Transformer (FoT)[Tworkowski et al., 2023], CodeLlama [Rozière et al., 2023] (CL) and YaRN Peng et al. [2023]. Table 7 presents the details about both standard and long-context pretraining. We use AdamW as an optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
Parameter/Model Size 270M 3B 7B 13B
Vocab Size 32000 32000 32000 32016
Embedding Size 1024 3200 4096 5120
Num Attention Layers 12 26 32 40
Num Attention Heads 8 32 32 40
Head Size 128 100 128 128
MLP Hidden Size 4096 8640 11008 13824
FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24] [12, 24, 36]
Each tuning experiment of 270M model was done using either a TPUv3-8 or TPUv4-8 machine and took around $10$ hours. Tuning a $3$B parameter model for $5.4$B tokens with FoT took $40$ hours on TPUv3-128. The $7$B and $13$B models were tuned on TPUv3-128.
Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model [Geng, 2023] and $13$B CodeLlama [Rozière et al., 2023] model. Subscript denotes that parameter was specific for a context extension method with FoT referring to [Tworkowski et al., 2023] and no-FoT to other methods (Naive, YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023]).
Stage Parameter/Model Size 270M 3B 7B 13B
Pretraining Context 2K 2K 2K 16K
Tokens 6.3B 1T 1T 2.5T
Long Context Tuning Context 32KFoT, 16Kno-FoT 32K 32K 32K
Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT 128
Start Learning Rate 5e-5 1.5e-5 1.5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6 1.5e-6 1.5e-6
Warmup Steps 250 1000 1000FoT, 200no-FoT 1000
Tokens 1B 5.4B 2B 2B
Appendix B Additional results for $270$M models
B.1 Short context evaluation
In Table 8, we asses the short context performance of $270$M models from Section 3.3 and compare them against their starting checkpoint.
Table 8: $2$K context perplexity evaluation of models from Section 3.3
Tune Dataset Method arXiv Code Code &
Haskell Python CUDA All arXiv
Starting Chkpt 10.80 13.40 6.00 5.00 7.17 7.40
Wikipedia SPLiCe BM25 10.64 13.28 6.00 5.00 7.16 7.38
SPLiCe Cont 10.60 13.24 5.99 4.99 7.14 7.36
Baseline 10.59 13.20 5.99 5.00 7.12 7.34
Stackexchange SPLiCe BM25 9.28 12.94 5.91 4.89 7.05 7.19
SPLiCe Cont 9.26 12.88 5.89 4.88 7.04 7.18
Baseline 9.24 13.00 5.91 4.90 7.07 7.21
C SPLiCe BM25 10.76 12.10 5.55 3.78 6.24 6.52
SPLiCe Cont 10.76 12.06 5.53 3.76 6.21 6.50
Baseline 10.73 11.96 5.49 3.68 6.16 6.44
SPLiCe Repo 10.78 12.10 5.56 3.79 6.24 6.52
C# SPLiCe BM25 10.86 12.74 5.73 4.67 6.72 6.98
SPLiCe Cont 10.89 12.72 5.72 4.65 6.70 6.96
Baseline 10.83 12.73 5.70 4.65 6.69 6.95
SPLiCe Repo 10.91 12.77 5.74 4.68 6.72 6.98
Python SPLiCe BM25 10.75 12.33 4.12 4.44 6.38 6.65
SPLiCe Cont 10.76 12.28 4.09 4.43 6.36 6.64
Baseline 10.72 12.19 3.98 4.40 6.33 6.61
SPLiCe Repo 10.78 12.33 4.12 4.45 6.38 6.66
B.2 Different context extension methods
We test SPLiCe with different context extension methods. In Table 9, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama [Rozière et al., 2023], and [Peng et al., 2023] RoPe [Su et al., 2021] adjustment method.
@@ B.2 Different context extension methods
RoPe Method arXiv Code Code &
scale Haskell Python CUDA All arXiv
Naive SPLiCe BM25 6.25 (.08) 4.84 (.19) 3.55 (.11) 2.84 (.12) 3.72 (.13) 3.88 (.12)
@@
B.3 Detailed results
In this section, we extend results presented in Section 3.3. Details about the number of tokens used for evaluation can be found in Table 19.